# 22-descent for Bloch–Kato Selmer groups and rational points on hyperelliptic curves I

### v3

@@ 1. Introduction
the effective computability was established by Baker [Bak69]. For rational points, effective computability is still an open problem, but the Chabauty–Coleman method is often practical for determining the set of rational points. However both of these methods have drawbacks which limit their applicability for practical computation. For Baker’s method, the limitation is that of determining integer points below within the height bound. For the Chabauty–Coleman method, one of the main limitations is the condition on the Mordell–Weil rank.
@@
In this article, we propose a naive approach more in the spirit of computing the rank of an abelian variety by a $2$-descent. Namely, given a $p$-adic representation $V$ of $\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q})$ with a Galois stable lattice $T$, one may bound the $\mathbb{Q}_{p}$-dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),V)$ by the $\mathbb{F}_{p}$-dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{p})$. Hence a first step is to be able to estimate the dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{p})$. We give such an estimate when $p=2$ and $V=\wedge^{2}V_{2}J$ where $J$ is the Jacobian of a hyperelliptic curve with a rational Weierstrass point, however the methods have a somewhat broader applicability. For example, if $J$ is as above then these methods give a (weaker) bound for the dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{2})$ when $T$ is any Galois representation generated by $V_{2}J$ (i.e. any summand of $V^{\otimes i}(j)$ for $i\geq 0$ and $j\in\mathbb{Z}$).
@@ Acknowledgements
I am grateful to Ed Schaefer for explaining $2$-Selmer groups of hyperelliptic curves to me, and to Michael Stoll for several detailed suggestions about an earlier version of this work, and corrections and comments on an earlier version of this paper. I am grateful to Lee Berry for many corrections to a previous version of this paper and useful discussions. This research was supported by a Royal Society University Research Fellowship.
@@ 2. Galois cohomology of the wedge square
Definition 1.
We define $K_{f}$ to be the étale $K$-algebra $K[x]/(f)$. We define $K_{f,2}$ to be the étale $K$-algebra $K[x,y,\frac{1}{x-y}]/(f(x),g(x))$. We define $\tau$ to be the involution of $K_{f,2}$ swapping the images of $x$ and $y$. We define $K_{f}^{(2)}$ to be the algebra of invariants $K_{f,2}^{\tau}$.
For an étale $K$-algebra $L$, with factorisation into fields $\prod_{i=1}^{n}L_{i}$, and a finite $\operatorname{Gal}(L^{\mathrm{sep}}|L)$-module $M$, we define $\mathrm{Ind}^{L}_{K}M:=\prod_{i=1}^{n}\mathrm{Ind}^{L_{i}}_{K}M$.
@@
Let $\mathrm{Nm}:\mathbb{F}_{2}[S^{(2)}]\to\mathbb{F}_{2}[S]$ denote the map sending $\{\alpha,\beta\}$ to $\alpha+\beta$. Let
@@
$K_{f}^{(2),\times}\otimes\mathbb{F}_{2}\to K_{f,2}^{\times}\otimes\mathbb{F}_{2}\times K_{f}\otimes\mathbb{F}_{2}$
@@ 3. The nonabelian (xT)(x-T)-map
For any root $\gamma$ of $f$ over an étale $K$-algebra $L$, we let $c_{\gamma}:=x(b)-\gamma$. If $z\in X(K)$ is a non–Weierstrass point with the property that zbz-b lies in $2\cdot J(K)$, then there is an element $u_{\gamma}$ of $L$ such that $c_{\gamma}u_{\gamma}^{2}=x(z)-\gamma$. For example, we can apply this construction to the étale algebra $K_{f,2}$, to get elements $c_{\alpha},u_{\alpha},c_{\beta}$ and $u_{\beta}$. The element $\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}$ will in fact be defined over the field $K_{f}^{(2)}$. This element of $K_{f}^{(2)}$ is related to $j_{J[4],b}(z)$ and $j_{\Gamma,b}(z)$ by the following ‘nonabelian (xT)(x-T) maps’.
@@
    $\mathrm{Nm}_{K_{f}^{(2)}|K_{f}}\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right).$
@@
    $2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right).$
@@
To descend $\mathbb{Z}/2\mathbb{Z}$ covers, we will make use of the following well-known construction.
@@
$\mathrm{Nm}_{K(\sqrt{d})|K}(x+y\sqrt{d})=c^{2},$
for $x,y,c\in K$, and $c$ nonzero. Then
$(x+y\sqrt{d})\equiv 2(x+c)$
@@
Then $K^{\mathrm{sep}}(X)(u_{\alpha})(\sqrt{g_{\alpha,0}})$ is an unramified $\mathbb{Z}/4\mathbb{Z}$-cover of $K^{\mathrm{sep}}(X)$. Isomorphism classes of lifts of $Y_{\alpha}\to X_{K_{f}}$ to an unramified $\mathbb{Z}/4\mathbb{Z}$-cover are in bijection with a choice of square root of $(\beta-\alpha)$ for all $\beta\neq\alpha$, modulo the equivalence relation
@@
(10) $w_{\alpha\beta}^{2}:=2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right).$
@@
$\mathrm{Nm}\left(\frac{u_{\alpha}-\sqrt{(\beta-\alpha)/c_{\alpha}}}{1-\sqrt{(\beta-\alpha)/c_{\alpha}}}\right)$ $=u_{\beta}$
@@
$\frac{2(\alpha-\beta)}{c_{\beta}}\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right).$
This is equal to $2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right)$ in $K^{\mathrm{sep}}(Y_{\alpha}\times_{X}Y_{\beta})^{\times}\otimes\mathbb{F}_{2}$. Finally, note that
$\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right)\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right)$
$=$ $-c_{\alpha}c_{\beta}\left(\frac{(u_{\beta}-1)(u_{\alpha}+1)}{(\alpha-\beta-c_{\beta}u_{\beta}-c_{\alpha}u_{\alpha})}\right).$
@@ 4. $2$-descent for Selmer schemes: local aspects away from 22
For any $G<G_{K,S}$, we obtain exact sequences
$0\to H^{1}(G,T)\otimes\mathbb{F}_{p}\to H^{1}(G,T\otimes\mathbb{F}_{p})\to H^{2}(G,T)[p]\to 0$
$0\to H^{0}(G,T)\otimes\mathbb{F}_{p}\to H^{0}(G,T\otimes\mathbb{F}_{p})\to H^{1}(G,T)[p]\to 0$
and an isomorphism $H^{1}(G,T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}\simeq H^{1}(G,V)$. We want to bound the rank of $H^{1}_{f}(G_{K,S},V)$ by controlling the rank of $H^{1}_{f}(G_{K,S},T)\otimes\mathbb{F}_{p}$. Our goal is to find subspaces of $H^{1}(G_{K,S},T\otimes\mathbb{F}_{p})$ which approximate $H^{1}_{f}(G_{K,S},T)\otimes\mathbb{F}_{p}$ and therefore whose $\mathbb{F}_{p}$-ranks give upper bounds on the $\mathbb{Q}_{p}$-rank of $H^{1}_{f}(G_{K,S},V)$.
@@ 4.1. Properties of $H^{1}_{f}(K_{v},T)$
Let $T$ be as above. Let $v$ be a place of $K$.
@@
    Suppose $v$ is prime to $p$ and $I_{v}$ acts unipotently on $T$. Then
@@
Part 1 is immediate from $H^{1}_{f}(K_{v},T\otimes\mathbb{F}_{p})$ being defined to be a subspace of the image of $H^{1}(K_{v},T)$. For part 2, by definition $H^{1}_{f}(K_{v},T)$ consists of the classes whose image in $H^{1}(I_{v},T)^{\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})}$ is torsion. Hence if the latter is torsion-free, then $H^{1}_{f}(K_{v},T)=H^{1}_{\mathrm{ur}}(K_{v},T)$. Since the image of $H^{1}_{\mathrm{ur}}(K_{v},T)$ in $H^{1}(K_{v},T\otimes\mathbb{F}_{p})$ is contained in $H^{1}_{\mathrm{ur}}(K_{v},T\otimes\mathbb{F}_{p})$, it is hence enough to show that $H^{1}(I_{v},T)^{\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})}$ is torsion-free when $I_{v}$ acts unipotently. Since $I_{v}$ acts unipotently, we have
@@
where $I_{v}^{(p)}$ is the maximal pro-p quotient of $I_{v}$. Since this is isomorphic, as a group with an outer action of $\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})$, to $\mathbb{Z}_{p}(1)$, we have an isomorphism of $\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})$-modules
$H^{1}(I_{v},M)\simeq(M/(\sigma-1))(-1),$
where $\sigma$ is a topological generator of $I_{v}^{(p)}$. Since $\sigma$ acts unipotently, $M/(\sigma-1)M$ is a torsion-free $\mathbb{Z}_{p}$-module, and hence so are the invariants (by any group action). ∎
@@
Lemma 18.
Suppose $X$ is a hyperelliptic curve with odd degree model $y^{2}=f(x)$ over a local field $K$ of residue characteristic different from $2$, and $X$ that has semistable reduction. Then
$H^{1}_{f}(K,J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K}^{\times}\otimes\mathbb{F}_{2})$
@@
$H^{1}_{f}(K,\wedge^{2}J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}^{(2)}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2})$
@@
We have
$H^{1}_{\mathrm{ur}}(K,J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K}^{\times}\otimes\mathbb{F}_{2})$
and
$H^{1}_{\mathrm{ur}}(K,\wedge^{2}J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}^{(2)}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2})$
By the semistable reduction theorem, $I_{v}$ acts unipotently on $T_{2}J$, and hence also on $\wedge^{2}T_{2}J$. Therefore the lemma follows from Proposition 17. ∎

### v4

@@ 1. Introduction
the effective computability was established by Baker [Bak69]. For rational points, effective computability is still an open problem, but the Chabauty–Coleman method is often practical for determining the set of rational points. However both of these methods have drawbacks which limit their applicability for practical computation. For Baker’s method, the limitation is that of determining integer points below the height bound. For the Chabauty–Coleman method, one of the main limitations is the condition on the Mordell–Weil rank.
@@
In this article, we propose a naive approach more in the spirit of computing the rank of an abelian variety via $2$-descent. Namely, given a $p$-adic representation $V$ of $\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q})$ with a Galois stable lattice $T$, one may bound the $\mathbb{Q}_{p}$-dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),V)$ by the $\mathbb{F}_{p}$-dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{p})$. Hence a first step is to be able to estimate the dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{p})$. We give such an estimate when $p=2$ and $V=\wedge^{2}V_{2}J$ where $J$ is the Jacobian of a hyperelliptic curve with a rational Weierstrass point, however the methods have a somewhat broader applicability. For example, if $J$ is as above then these methods give a (weaker) bound for the dimension of $H^{1}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{2})$ when $T$ is any Galois representation generated by $V_{2}J$ (i.e. any summand of $V^{\otimes i}(j)$ for $i\geq 0$ and $j\in\mathbb{Z}$).
@@ Acknowledgements
I would like to thank the anonymous referees for helpful corrections and suggestions for changes to the structure of the paper. I am grateful to Ed Schaefer for explaining $2$-Selmer groups of hyperelliptic curves to me, and to Michael Stoll for several detailed suggestions about an earlier version of this work, and corrections and comments on an earlier version of this paper. I am grateful to Lee Berry for many corrections to a previous version of this paper and useful discussions. This research was supported by a Royal Society University Research Fellowship.
@@ 2. Galois cohomology of the wedge square
Recall from the introduction the étale algebras $K_{f},K_{f,2}$ and $K_{f}^{(2)}$. For an étale $K$-algebra $L$, with factorisation into fields $\prod_{i=1}^{n}L_{i}$, and a finite abelian group with an action of $\operatorname{Gal}(L_{i})$ for each $i$, we define $\mathrm{Ind}^{L}_{K}M:=\oplus_{i=1}^{n}\mathrm{Ind}^{L_{i}}_{K}M$.
@@
Let $\mathrm{Nm}:\mathbb{F}_{2}[S^{(2)}]\to\mathbb{F}_{2}[S]$ denote the map sending $\{\alpha,\beta\}]$ to $\alpha]+[\beta]$. Let
@@
$K_{f}^{(2),\times}\otimes\mathbb{F}_{2}\to K_{f,2}^{\times}\otimes\mathbb{F}_{2}\to K_{f}\otimes\mathbb{F}_{2}$
@@ 3. The nonabelian (xT)(x-T)-map
For any root $\gamma$ of $f$ over an étale $K$-algebra $L$, we let $c_{\gamma}:=x(b)-\gamma$. If $z\in X(K)$ is a non–Weierstrass point with the property that zbz-b lies in $2\cdot J(K)$, then there is an element $u_{\gamma}$ of $L$ such that $c_{\gamma}u_{\gamma}^{2}=x(z)-\gamma$. For example, we can apply this construction to the étale algebra $K_{f,2}$, to get elements $c_{\alpha},u_{\alpha},c_{\beta}$ and $u_{\beta}$. The element $\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}$ will in fact be defined over the field $K_{f}^{(2)}$. This element of $K_{f}^{(2)}$ is related to $j_{J[4],b}(z)$ and $j_{\Gamma,b}(z)$ by the following ‘nonabelian (xT)(x-T) maps’.
@@
    $\mathrm{Nm}_{K_{f}^{(2)}|K_{f}}\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}\right).$
@@
    $2\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}\right).$
@@
To descend $\mathbb{Z}/2\mathbb{Z}$ covers, we will make use of the following well-known calculation.
@@
$\mathrm{Nm}_{K(\sqrt{d})|K}(x+y\sqrt{d})=z^{2},$
for $x,y,z\in K$, and $z$ nonzero. Then
$x+y\sqrt{d}=\frac{(x+z+y\sqrt{d})^{2}}{2(x+z)}.$
In particular,
$(x+y\sqrt{d})\equiv 2(x+z)$
@@
Then $K^{\mathrm{sep}}(X)(u_{\alpha})(\sqrt{g_{\alpha,0}})$ is a $\mathbb{Z}/4\mathbb{Z}$-extension of $K^{\mathrm{sep}}(X)$ extending to an étale cover of $X_{K^{\mathrm{sep}}}$. Isomorphism classes of lifts of $Y_{\alpha}\to X_{K_{f}}$ to an unramified $\mathbb{Z}/4\mathbb{Z}$-cover are in bijection with a choice of square root of $(\beta-\alpha)$ for all $\beta\neq\alpha$, modulo the equivalence relation
@@
(10) $w_{\alpha\beta}^{2}:=2\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}\right).$
@@
$\mathrm{Nm}\left(\frac{u_{\alpha}-\sqrt{(\beta-\alpha)/c_{\alpha}}}{1-\sqrt{(\beta-\alpha)/c_{\alpha}}}\right)$ $=u_{\beta}^{2}$
@@
$\frac{2(\alpha-\beta)}{c_{\beta}}\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\alpha-\beta}\right).$
This is equal to $2\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\alpha-\beta}\right)$ in $K^{\mathrm{sep}}(Y_{\alpha}\times_{X}Y_{\beta})^{\times}\otimes\mathbb{F}_{2}$. Finally, note that
$\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\alpha-\beta}\right)\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}\right)$
$=$ $-c_{\alpha}c_{\beta}\left(\frac{(u_{\beta}-1)(u_{\alpha}+1)}{(\alpha-\beta-c_{\beta}u_{\beta}-c_{\alpha}u_{\alpha})}\right)^{2}.$
@@ 4. $2$-descent for Selmer schemes: local aspects away from 22
For any $G<G_{K,S}$ and $i\geq 0$, we obtain an exact sequence
(15) $0\to H^{i}(G,T)\otimes\mathbb{F}_{p}\to H^{i}(G,T\otimes\mathbb{F}_{p})\to H^{i+1}(G,T)[p]\to 0$
and an isomorphism $H^{i}(G,T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}\simeq H^{i}(G,V)$. We want to bound the rank of $H^{1}_{f}(G_{K,S},V)$ by controlling the rank of $H^{1}_{f}(G_{K,S},T)\otimes\mathbb{F}_{p}$. Our goal is to find subspaces of $H^{1}(G_{K,S},T\otimes\mathbb{F}_{p})$ which approximate $H^{1}_{f}(G_{K,S},T)\otimes\mathbb{F}_{p}$ and therefore whose $\mathbb{F}_{p}$-ranks give upper bounds on the $\mathbb{Q}_{p}$-rank of $H^{1}_{f}(G_{K,S},V)$.
@@ 4.1. Properties of $H^{1}_{f}(K_{v},T)$
Let $T$ be as above. Let $v$ be a place of $K$, $k$ the residue field of the ring of integers of $K_{v}$, and $\sigma$ a topological generator for the tame inertia subgroup of the Galois group of $K_{v}$.
@@
    Suppose $v$ is prime to $p$, $I_{v}$ acts unipotently on $T$, and $T/(\sigma-1)T$ is torsion-free. Then
@@
Part 1 is immediate from (15). For part 2, by definition $H^{1}_{f}(K_{v},T)$ consists of the classes whose image in $H^{1}(I_{v},T)^{\operatorname{Gal}(\overline{k}_{v}|k_{v})}$ is torsion. Hence if the latter is torsion-free, then $H^{1}_{f}(K_{v},T)=H^{1}_{\mathrm{ur}}(K_{v},T)$. Since the image of $H^{1}_{\mathrm{ur}}(K_{v},T)$ in $H^{1}(K_{v},T\otimes\mathbb{F}_{p})$ is contained in $H^{1}_{\mathrm{ur}}(K_{v},T\otimes\mathbb{F}_{p})$, it is hence enough to show that $H^{1}(I_{v},T)^{\operatorname{Gal}(\overline{k}_{v}|k_{v})}$ is torsion-free when $I_{v}$ acts unipotently. Since $I_{v}$ acts unipotently, we have
@@
where $I_{v}^{(p)}$ is the maximal pro-p quotient of $I_{v}$. Hence we have an isomorphism of $\mathbb{Z}_{p}$-modules
$H^{1}(I_{v},M)\simeq M/(\sigma-1)M$
where $\sigma$ is a topological generator of $I_{v}^{(p)}$. Since $M/(\sigma-1)M$ is a torsion-free $\mathbb{Z}_{p}$-module, so are the invariants (by any group action). ∎
@@
Recall that, given an oriented graph $\Gamma$ with vertex set $V$ and edge set $E$, we have maps
$\delta:\mathbb{Z}[E]\to\mathbb{Z}[V]$
$e\mapsto s(e)-t(e).$
@@
$\epsilon:\mathbb{Z}[V]\to\mathbb{Z}[E]$
$v\mapsto\sum_{s(e)=v}e-\sum_{t(e)=v}v$
We define
$H_{1}(\Gamma,\mathbb{Z}):=\mathrm{Ker}(\delta)$
$H^{1}(\Gamma,\mathbb{Z}):=\mathrm{Coker}(\epsilon)$
and define the monodromy map
$N_{\Gamma}:H_{1}(\Gamma,\mathbb{Z})\to H^{1}(\Gamma,\mathbb{Z})$
to be the composite of the inclusion of $H_{1}(\Gamma,\mathbb{Z})$ into $\mathbb{Z}[E]$ with the quotient map $\mathbb{Z}[E]\to H^{1}(\Gamma,\mathbb{Z})$. Although the definitions above are phrased in terms of an oriented graph, for two different choices of an orientation on a graph $\Gamma$ (i.e. choices of an ordering of the endpoints of each edge) the homology and cohomology groups above are isomorphic in a manner compatible with the monodromy map. For this reason we will view $H_{1}(\Gamma,\mathbb{Z}),H^{1}(\Gamma,\mathbb{Z})$ and $N_{\Gamma}$ as having input a graph $\Gamma$.
Now let $K_{v}$ and $\sigma$ be as above. Let $\mathcal{O}_{v}$ denote the ring of integers of $K_{v}$ and $k_{v}$ the residue field. Let $X/K_{v}$ be a smooth projective geometrically irreducible curve with strict semistable reduction and $X/\mathcal{O}_{v}$ be a minimal regular model. Let $\Gamma$ denote the dual graph of $X_{k_{v}}$. Let $J$ be the Jacobian of $X$ and let $p$ be a prime which is a unit in $\mathcal{O}_{v}$. Grothendieck’s Picard–Lefschetz formula describes the action of $\sigma$ on $T_{p}J$ in terms of $N_{\Gamma}$.
Theorem 3.
[Picard–Lefschetz formula, [SGA72]]
$T_{p}J$ admits a Galois stable filtration
$T_{p}J\supset W_{1}T_{p}J\supset W_{2}T_{p}J.$
The map
$(\sigma-1):T_{p}J\to T_{p}J$
has kernel equal to $W_{1}T_{p}J$ and image contained in $W_{2}T_{p}J$. We have Galois equivariant isomorphisms
$T_{p}J/W_{1}T_{p}J\simeq H_{1}(\Gamma,\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Z}_{p}$
$W_{2}T_{p}J\simeq H^{1}(\Gamma,\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Z}_{p}(1)$
giving a commutative diagram
[…]
where $\alpha\in\mathbb{Z}_{p}(1)$ is an element corresponding to $\sigma$.
@@
For the convenience of the reader we recall how to extract this statement from [SGA72, Exposé VIII Theorem 12.5]. First let $\mathcal{A}$ denote the Neron model of $J$ over $\mathcal{O}_{v}$. Let $\mathcal{A}^{o}_{k_{v}}\subset\mathcal{A}_{k_{v}}$ denote the connected component of the identity of the special fibre, and let $T\subset\mathcal{A}^{o}_{k_{v}}$ denote the toric part. Then the filtration is given by
$W_{1}T_{p}J=T_{p}\mathcal{A}^{o}_{k_{v}}$
$W_{2}T_{p}J=T_{p}T.$
In loc. cit., it is proved that two monodromy pairings coincide. The first is the bilinear pairing [SGA72, 12.4.5]
$\Delta:H_{1}(\Gamma,\mathbb{Z})\times H_{1}(\Gamma,\mathbb{Z})\to\mathbb{Z}$
obtained by restriction from the bilinear pairing
$\Delta:\mathbb{Z}[E]\times\mathbb{Z}[E]\to\mathbb{Z}$
which sends $(e_{1},e_{2})$ to $1$ if $e_{1}=e_{2}$ and 0 otherwise.
On the other hand the Weil pairing defines an isomorphism $\mathrm{Hom}(W_{2}T_{p}J,\mathbb{Z}_{p}(1))\simeq T_{p}J/W_{1}T_{p}J$, and hence $\sigma-1$ defines a map
$T_{p}J/W_{1}T_{p}J\to\mathrm{Hom}(T_{p}J/W_{1}T_{p}J,\mathbb{Z}_{p}(1)),$
or dually a pairing
$T_{p}J/W_{1}T_{p}J\times T_{p}J/W_{1}T_{p}J\to\mathbb{Z}_{p}(1).$
The statement of the Picard–Lefschetz theorem in loc. cit. implies that this pairing is identified with
$\Delta\otimes\alpha:H_{1}(\Gamma,\mathbb{Z})\otimes\mathbb{Z}_{p}\times H_{1}(\Gamma,\mathbb{Z})\otimes\mathbb{Z}_{p}\to\mathbb{Z}_{p}(1)$
via an isomorphism $H_{1}(\Gamma,\mathbb{Z})\otimes\mathbb{Z}_{p}\simeq T_{p}J/W_{1}T_{p}J$. Dualising gives the statement of the theorem above. ∎
Corollary 1.
  1. (1)
    Suppose $X/K_{v}$ is a smooth projective geometrically irreducible curve with regular model whose special fibre is a curve of genus g1g-1 with two points identified. Let $T=\wedge^{2}T_{p}J$. Then $I_{v}$ acts unipotently on $T$, and $T/(\sigma-1)T$ is torsion free.
  2. (2)
    If $v$ lies above an odd prime, $f\in\mathcal{O}_{v}[x]$ is a polynomial of degree $\geq 5$ which is separable over $K_{v}$ with leading coefficient a $v$-adic unit, and $\mathrm{val}_{v}(\Delta(f))=1$, then $H^{1}_{f}(K_{v},\wedge^{2}J[2])=H^{1}_{\mathrm{ur}}(K_{v},\wedge^{2}J[2])$.
Proof.
For part (1), let $\Gamma$ denote the dual graph of a minimal regular model of $X$ over $\mathbb{Z}_{\ell}$. Then the monodromy map
$N:H_{1}(\Gamma,\mathbb{Z})\to H^{1}(\Gamma,\mathbb{Z})$
is an isomorphism. This implies $(\sigma-1)T_{p}J$ is torsion-free, and hence the same is true of $(\sigma-1)\wedge^{2}T_{p}J$. For part (2), note that the condition on the discriminant implies that $y^{2}=f(x)$ defines a minimal regular model as in part (1). Hence the conclusion follows from part (1), Theorem 3 and Lemma 17. ∎